# Quantum geometric tensor determines the i.i.d. conversion rate in the resource theory of asymmetry for any compact Lie group

### v2

@@ Acknowledgements.
The authors would like to thank Yui Kuramochi for valuable comments on an early draft of this manuscript, and also Iman Marvian for the valuable discussion. K.Y. acknowledges support from JSPS KAKENHI Grant No. JP24KJ0085. Y.M. is supported by JSPS KAKENHI Grant No. JP23KJ0421. T.S. acknowledges support from JST Moonshot R&D Grant No. JPMJMS2061, JST CREST Grant No. JPMJCR23I4, and MEXT Q-LEAP Grant No. JPMXS0120319794, Japan. H.T. was supported by MEXT KAKENHI Grant-in-Aid for Transformative Research Areas B “Quantum Energy Innovation” Grant Numbers 24H00830 and 24H00831, and JST PRESTO No. JPMJPR2014, JST MOONSHOT No. JPMJMS2061.
@@ A.1 Differentiability of representation
For projective unitary representations $U$ and $U^{\prime}$ that are continuous, (non-projective) unitary representations $\tilde{U}$ and $\tilde{U}^{\prime}$ in Eq. (78) are continuous and hence differentiable. Therefore, the conversion rate among pure state is obtained from Eq. (79) since its right-hand side can be calculated from Theorem 1.
@@
where $X_{\mu}\coloneqq\left.-\mathrm{i}\frac{\partial}{\partial\lambda^{\mu}}U(g(\bm{\lambda}))\right|_{\bm{\lambda}=\bm{0}}$. Similarly, the QGT for $\mathcal{U}_{g}(\psi)=\ket{\psi(g)}\bra{\psi(g)}$ |, where |$\ket{\psi(g)}\coloneqq U(g)\ket{\psi}$, is given by
$\mathcal{Q}^{\mathcal{U}_{g}(\psi)}_{\mu\nu}$ $=\braket{\psi(g)}{X_{\mu}(I-\ket{\psi(g)}\bra{\psi(g)})X_{\nu}}{\psi(g)}$ (83)
@@ A.3 Conversion rates for projective unitary representations and (non-projective) unitary representations
Let $U$ be a projective unitary projective representation of a Lie group $G$ on a $d$-dimensional Hilbert space $\mathcal{H}$. Define a (non-projective) unitary representation $\tilde{U}$ by Eq. (78). Then it holds
@@
$\frac{1}{d}\mathcal{Q}_{\mu\nu}^{\psi^{\otimes d}}$ $=\frac{1}{d}\sum_{i=1}^{d}\left(\braket{\partial_{\mu}U(g(\bm{\lambda}))\psi}{\partial_{\nu}U(g(\bm{\lambda}))\psi}-\braket{\partial_{\mu}U(g(\bm{\lambda}))\psi}{U(g(\bm{\lambda}))\psi}\braket{U(g(\bm{\lambda}))\psi}{\partial_{\nu}U(g(\bm{\lambda}))\psi}\right)\biggl{|}_{\bm{\lambda}=\bm{0}}$
@@ A.4 Symmetry subgroup and the conversion rate
We here prove that $R(\psi\to\phi)=0$ unless $\mathrm{Sym}_{G}(\psi)\subset\mathrm{Sym}_{G}(\phi)$. In fact, we can prove this fact not only for pure states but also for generic states:
@@ B.1 Proof of Lemma 3
In order to relate the above notation with ours, let us now consider a particular probability distribution $p_{q}\coloneqq(1-q)\delta_{1}+q\delta_{0}$ characterized by $q\in(0,1)$, where $\delta_{0}$ and $\delta_{1}$ denote the Dirac measure concentrated on 00 and $1$, respectively. In this case, we have
@@
Let $\rho=\sum_{i=1}^{d}p_{i}\ket{i}\bra{i}$ | be the eigenvalue decomposition of $\rho$. For an operator $A$ defined by
$A\coloneqq\sum_{k,l;m_{f_{q}}(p_{k},p_{l})>0}\frac{1}{m_{f_{q}}(p_{k},p_{l})}\ket{l}\braket{l}{\mathrm{i}[\rho,B]}{k}\bra{k},$ (110)
@@
$E_{p_{q},\rho}(A)$ $=\sum_{k,l;(1-q)p_{l}+qp_{k}>0}\frac{(1-q)p_{l}}{(1-q)p_{l}+qp_{k}}\ket{l}\braket{l}{\mathrm{i}[\rho,B]}{k}\bra{k}+\sum_{k,l;(1-q)p_{l}+qp_{k}>0}\frac{qp_{k}}{(1-q)p_{l}+qp_{k}}\ket{l}\braket{l}{\mathrm{i}[\rho,B]}{k}\bra{k}$ | (111)
$=\sum_{k,l;(1-q)p_{l}+qp_{k}>0}\ket{l}\braket{l}{\mathrm{i}[\rho,B]}{k}\bra{k}$ | (112)
$=\mathrm{i}[\rho,B]-\sum_{k,l;p_{l}=p_{k}=0}\ket{l}\braket{l}{\mathrm{i}[\rho,B]}{k}\bra{k}$ | (113)
@@
$\left(\|\mathrm{i}[\rho,B]\|_{p_{q},\rho}^{(m)}\right)^{2}$ $=\mathrm{Tr}\left(\left(\sum_{k,l;m_{f_{q}}(p_{k},p_{l})>0}\frac{1}{m_{f_{q}}(p_{k},p_{l})}\ket{l}\braket{l}{\mathrm{i}[\rho,B]}{k}\bra{k}\right)^{\dagger}(\mathrm{i}[\rho,B])\right)$ (115)
@@ B.2 Proof of the equivalence between two alternative definitions of asymptotic convertibility
(ii)$\implies$(i): Any sequence of $G$-covariant channels $\{\mathcal{E}_{N}\}_{N}$ such that $\lim_{N\to\infty}T({\mathcal{E}}_{N}(\rho_{N}),\sigma_{N})=0$ provides an example of the channels satisfying condition (i).
(i)$\implies$(ii): Fix a sequence of positive numbers $\{\epsilon_{k}\}_{k}$ such that $\lim_{k\to\infty}\epsilon_{k}=0$. Assume the condition (i) is true. Then for each $\epsilon_{k}>0$, $\exists N_{0}(\epsilon_{k})>0$ such that $\forall N>N_{0}(\epsilon_{k})$, $\exists{\mathcal{E}}_{N,\epsilon_{k}}:G\text{-cov.}\ {\rm s.t.}\ T({\mathcal{E}}_{N}(\rho_{N}),\sigma_{N})\leq\epsilon_{k}$. Without loss of generality, we can assume $N_{0}(\epsilon_{k+1})>N_{0}(\epsilon_{k})$. Let us define $G$-covariant channels $\{\mathcal{E}_{N}\}_{N}$ by
@@ B.3 Proof of Lemma 4
Let $\sigma$ be an arbitrary state and $\phi$ be a pure state. Suppose that their infidelity by $\delta\coloneqq 1-\braket{\phi}{\sigma}{\phi}$ satisfies $\delta<1/2$. For the eigenvector $\Phi$ with the largest eigenvalue of $\sigma$, it holds
$|\braket{\phi}{\Phi}|^{2}\geq 1-2\delta.$ (118)
@@
$\braket{\mathrm{i}[\sigma,O],\mathrm{i}[\sigma,O]}_{f,\sigma}\geq\frac{1}{f\left(\frac{\delta}{1-\delta}\right)}(1-2\delta)^{2}\Braket{\Phi}{O(I-\Phi)O^{\dagger}}{\Phi}.$ (119)
@@
where $\lambda_{0}\coloneqq p$ and |$\ket{0}\coloneqq\ket{\Phi}$. Since $p$ is assumed to be the largest eigenvalue, we have
$1-\delta=p|\braket{\phi}{\Phi}|^{2}+\sum_{k=1}^{d-1}\lambda_{k}|\braket{\phi}{k}|^{2}\leq p\sum_{k=0}^{d-1}|\braket{\phi}{k}|^{2}=p.$ (121)
@@
$1-\delta$ $=\braket{\phi}{\sigma}{\phi}$ (122)
$=p|\braket{\phi}{\Phi}|^{2}+\sum_{k=1}^{d-1}\lambda_{k}|\braket{\phi}{k}|^{2}$ (123)
$\leq p|\braket{\phi}{\Phi}|^{2}+(1-p)\sum_{k=1}^{d-1}|\braket{\phi}{k}|^{2}$ (124)
$\leq p|\braket{\phi}{\Phi}|^{2}+\delta(1-|\braket{\phi}{\Phi}|^{2}),$ (125)
@@
$|\braket{\phi}{\Phi}|^{2}\geq\frac{1-2\delta}{p-\delta}\geq 1-2\delta,$ (126)
@@
$\braket{\mathrm{i}[\sigma,O],\mathrm{i}[\sigma,O]}_{f,\sigma}$ $=\sum_{k,l;\lambda_{l}f(\lambda_{k}/\lambda_{l})>0}\frac{1}{\lambda_{l}f(\lambda_{k}/\lambda_{l})}\braket{k}{[\sigma,O^{\dagger}]}{l}\braket{l}{[O,\sigma]}{k}$ (127)
$=\sum_{k,l;\lambda_{l}f(\lambda_{k}/\lambda_{l})>0}\frac{(\lambda_{l}-\lambda_{k})^{2}}{\lambda_{l}f(\lambda_{k}/\lambda_{l})}|\braket{l}{O}{k}|^{2}+\sum_{k,l;\lambda_{l}f(\lambda_{k}/\lambda_{l})>0,l\neq 0}\frac{(\lambda_{l}-\lambda_{k})^{2}}{\lambda_{l}f(\lambda_{k}/\lambda_{l})}|\braket{l}{O}{k}|^{2}$ (128)
$\geq\sum_{k;pf(\lambda_{k}/p)>0}\frac{(p-\lambda_{k})^{2}}{pf(\lambda_{k}/p)}|\braket{\Phi}{O}{k}|^{2}$ (129)
$=\sum_{k=1}^{d-1}\frac{(p-\lambda_{k})^{2}}{pf(\lambda_{k}/p)}|\braket{\Phi}{O}{k}|^{2},$ (130)
@@
$\braket{\mathrm{i}[\sigma,O],\mathrm{i}[\sigma,O]}_{f,\sigma}$ $\geq\sum_{k=1}^{d-1}\frac{(p-\lambda_{k})^{2}}{pf(\lambda_{k}/p)}|\braket{\Phi}{O}{k}|^{2}$ (131)
$\geq\frac{1}{f\left(\frac{\delta}{1-\delta}\right)}(1-2\delta)^{2}\sum_{k=1}^{d-1}|\braket{\Phi}{O}{k}|^{2}$ (132)
$=\frac{1}{f\left(\frac{\delta}{1-\delta}\right)}(1-2\delta)^{2}\braket{\Phi}{O(I-\Phi)O^{\dagger}}{\Phi}.$ (133)
@@
$\Braket{\Phi_{N}}{O_{N}(I-\Phi_{N})O_{N}^{\dagger}}{\Phi_{N}}$ $=V(\Phi_{N},A_{N})+V(\Phi_{N},B_{N})-\mathrm{i}\braket{\Phi_{N}}{[A_{N},B_{N}]}{\Phi_{N}},$ (139)
[…]
@@
$-\mathrm{i}\braket{\Phi_{N}}{[A_{N},B_{N}]}{\Phi_{N}}$ $\geq-\mathrm{i}\braket{\phi^{\otimes N}}{[A_{N},B_{N}]}{\phi^{\otimes N}}-N\|\mathrm{i}[A,B]\|_{\infty}\|\Phi_{N}-\phi^{\otimes N}\|_{1}.$ (146)
@@
Let $A$ be an arbitrary Hermitian operator on a finite-dimensional Hilbert space $\mathcal{H}$. We denote their i.i.d. extensions by $A_{N}$ defined in Eq. (137). Fix a pure state |$\ket{\phi}\in\mathcal{H}$ and a sufficiently small parameter $\epsilon$. There exists a function $h_{1}(\epsilon)$ independent of $N$ satisfying $\lim_{\epsilon\to 0}h_{1}(\epsilon)=0$ such that for any sequence of pure states $\{\Phi_{N}\}_{N}$ satisfying $\limsup_{N\to\infty}T(\Phi_{N},\phi^{\otimes N})<\epsilon$, it holds
@@
[…]
$Q_{N}$ $\coloneqq(A_{N}-\braket{\Phi_{N}}{A_{N}}{\Phi_{N}}I)^{2}=(A_{N}-\braket{\phi^{\otimes N}}{A_{N}}{\phi^{\otimes N}}I+\Delta\mu_{N}I)^{2}.$ (149)
@@
$\braket{\phi^{\otimes N}}{(A_{N}-\braket{\Phi_{N}}{A_{N}|\Phi_{N}}I)^{2}|\phi^{\otimes N}}{=}\braket{\phi^{\otimes N}}{(A_{N}-\braket{\phi^{\otimes N}}{A_{N}|\phi^{\otimes N}}I)^{2}|\phi^{\otimes N}}{+}|\Delta\mu_{N}|^{2},$ (150)
@@
Here, we derive an upper bound of $\mathrm{Tr}\left(Q_{N}(\phi^{\otimes N}-\Phi_{N})\right)$. Let us denote the eigenvalue decomposition of $A$ by $A=\sum_{j=1}^{d}a_{j}\ket{j}\bra{j}$ |. Its i.i.d. extension is written as
$A_{N}=\sum_{j_{1}=1}^{d-1}\cdots\sum_{j_{N}=1}^{d-1}s(\bm{j})\ket{\bm{j}}\bra{\bm{j}},\quad\ket{\bm{j}}\coloneqq\ket{j_{1}\otimes j_{2}\otimes\cdots\otimes j_{N}},\quad s(\bm{j})\coloneqq\sum_{n=1}^{N}a_{j_{n}},$ (152)
@@
Since $Q_{N}$ is diagonal in {|$\{\ket{\bm{j}}\}$ basis, we can also decompose $Q_{N}$ as follows:
@@
$Q_{N}^{\mathrm{(core)}}(x)$ $\coloneqq\sum_{\bm{j}\in\mathcal{A}^{(\mathrm{core})}(x)}(s(\bm{j})-N\mu+\Delta\mu_{N})^{2}\ket{\bm{j}}\bra{\bm{j}},$ (155)
$Q_{N}^{\mathrm{(tail)}}(x)$ $\coloneqq\sum_{\bm{j}\in\mathcal{A}^{(\mathrm{tail})}(x)}(s(\bm{j})-N\mu+\Delta\mu_{N})^{2}\ket{\bm{j}}\bra{\bm{j}}.$ | . (156)
@@
Let $p$ denote a probability distribution defined by $p(j)\coloneqq|\braket{j}{\phi}|^{2}$. We denote its i.i.d. extension by
@@
For sufficiently small $\epsilon>0$ satisfying $01-2\epsilon-g_{0}(\epsilon)>$ 0, we have
@@
holds for all sufficiently large $N$, as long as $\epsilon>0$ is sufficiently small so that $01-2\epsilon-g_{0}(\epsilon)>$ 0. ∎
From Lemmas 22 and 23, we obtain the following corollary:
@@
Let $O$ be an arbitrary linear operator on a finite-dimensional Hilbert space $\mathcal{H}$. We denote by $O_{N}$ its i.i.d. extension defined in Eq. (134). Fix a pure state |$\ket{\phi}\in\mathcal{H}$ and a sufficiently small parameter $\epsilon$. Then there exists a function $h_{2}(\epsilon)$ independent of $N$ satisfying $\lim_{\epsilon\to 0}h_{2}(\epsilon)=0$ such that for any sequence of pure states $\{\Phi_{N}\}_{N}$ satisfying $\limsup_{N\to\infty}T(\Phi_{N},\phi^{\otimes N})<\epsilon$, it holds
$\Braket{\Phi_{N}}{O_{N}(I-\Phi_{N})O_{N}^{\dagger}}{\Phi_{N}}\geq\braket{\phi^{\otimes N}}{O_{N}(I-\phi^{\otimes N})O_{N}^{\dagger}}{\phi^{\otimes N}}-Nh_{2}(\epsilon)$ (185)
@@
$\Braket{\Phi_{N}}{O_{N}(I-\Phi_{N})O_{N}^{\dagger}}{\Phi_{N}}\geq\braket{\phi^{\otimes N}}{O_{N}(I-\phi^{\otimes N})O_{N}^{\dagger}}{\phi^{\otimes N}}-Nh_{2}(\epsilon)$ (187)
@@
$\braket{\mathrm{i}[\sigma_{N},O],\mathrm{i}[\sigma_{N},O]}_{f,\sigma_{N}}\geq\frac{1}{f\left(\frac{\delta_{N}}{1-\delta_{N}}\right)}(1-2\delta_{N})^{2}\Braket{\Phi_{N}}{O_{N}(I-\Phi)O_{N}^{\dagger}}{\Phi_{N}},$ (190)
where $\delta_{N}\coloneqq 1-\braket{\phi^{\otimes N}}{\sigma_{N}}{\phi^{\otimes N}}$ denotes the infidelity. For a sufficiently small $\epsilon>0$, from Lemma 24, we further find
$\braket{\mathrm{i}[\sigma_{N},O],\mathrm{i}[\sigma_{N},O]}_{f,\sigma_{N}}\geq\frac{1}{f\left(\frac{\delta_{N}}{1-\delta_{N}}\right)}(1-2\delta_{N})^{2}\left(N\braket{\phi}{O(I-\phi)O^{\dagger}}{\phi}-Nh_{2}(\epsilon)\right)$ (191)
@@
$f(0)\braket{\mathrm{i}[\sigma_{N},O],\mathrm{i}[\sigma_{N},O]}_{f,\sigma_{N}}\geq\frac{f(0)}{f\left(\frac{\delta_{N}}{1-\delta_{N}}\right)}(1-2\delta_{N})^{2}N\braket{\phi}{O(I-\phi)O^{\dagger}}{\phi}-Nh(\epsilon),$ (192)
@@
$f(0)\braket{\mathrm{i}[\sigma_{N},O],\mathrm{i}[\sigma_{N},O]}_{f,\sigma_{N}}\geq N\braket{\phi}{O(I-\phi)O^{\dagger}}{\phi}-Nh(\epsilon)+N\braket{\phi}{O(I-\phi)O^{\dagger}}{\phi}\left(\frac{f(0)}{f\left(\frac{\delta_{N}}{1-\delta_{N}}\right)}(1-2\delta_{N})^{2}-1\right).$ (194)

### v3

@@ Acknowledgements.
The authors would like to thank Yui Kuramochi for valuable comments on an early draft of this manuscript, and also Iman Marvian for the valuable discussion. The authors would also like to thank Mark M. Wilde and Takahiro Sagawa for separately pointing out that the conversion formula can also be rewritten using the max-relative entropy. K.Y. acknowledges support from JSPS KAKENHI Grant No. JP24KJ0085. Y.M. is supported by JSPS KAKENHI Grant No. JP23KJ0421. T.S. acknowledges support from JST Moonshot R&D Grant No. JPMJMS2061, JST CREST Grant No. JPMJCR23I4, and MEXT Q-LEAP Grant No. JPMXS0120319794, Japan. H.T. was supported by MEXT KAKENHI Grant-in-Aid for Transformative Research Areas B “Quantum Energy Innovation” Grant Numbers 24H00830 and 24H00831, and JST PRESTO No. JPMJPR2014, JST MOONSHOT No. JPMJMS2061.
@@ A.1 Differentiability of representation
For projective unitary representations $U$ and $U^{\prime}$ that are continuous, (non-projective) unitary representations $\tilde{U}$ and $\tilde{U}^{\prime}$ in Eq. (122) are continuous and hence differentiable. Therefore, the conversion rate between pure states is obtained from Eq. (123) since its right-hand side can be calculated from Theorem 1.
@@ A.3 Conversion rates for projective unitary representations and (non-projective) unitary representations
Let $U$ be a projective unitary representation of a Lie group $G$ on a $d$-dimensional Hilbert space $\mathcal{H}$. Define a (non-projective) unitary representation $\tilde{U}$ by Eq. (122). Then it holds
@@
$\frac{1}{d}\mathcal{Q}_{\mu\nu}^{\psi^{\otimes d}}$ $=\frac{1}{d}\sum_{i=1}^{d}\left(\braket{\partial_{\mu}U(g(\bm{\lambda}))\psi|\partial_{\nu}U(g(\bm{\lambda}))\psi}-\braket{\partial_{\mu}U(g(\bm{\lambda}))\psi|U(g(\bm{\lambda}))\psi}\braket{U(g(\bm{\lambda}))\psi|\partial_{\nu}U(g(\bm{\lambda}))\psi}\right)\biggl{|}_{\bm{\lambda}=\bm{0}}$
@@ A.4 Symmetry subgroup and the conversion rate
We here prove that $R(\psi\to\phi)=0$ unless $\mathrm{Sym}_{G}(\psi)\subset\mathrm{Sym}_{G}(\phi)$. In fact, we can prove this fact not only for pure states but also for any states:
@@ B.1 Proof of Lemma 3
In order to relate the above notation with ours, let us now consider a particular probability distribution $p_{q}\coloneqq(1-q)\delta_{1}+q\delta_{0}$ characterized by $q\in(0,1)$, where $\delta_{0}$ and $\delta_{1}$ denote the Dirac measure concentrated on 0 and $1$, respectively. In this case, we have
@@
Let $|\rho=\sum_{i=1}^{d}p_{i}\ket{i}\bra{i}$ be the eigenvalue decomposition of $\rho$. For an operator $A$ defined by
$|,\displaystyle A\coloneqq\sum_{k,l;m_{f_{q}}(p_{k},p_{l})>0}\frac{1}{m_{f_{q}}(p_{k},p_{l})}\ket{l}\braket{l|\mathrm{i}[\rho,B]|k}\bra{k}$, (154)
@@
[…]
$|\displaystyle=\sum_{k,l;(1-q)p_{l}+qp_{k}>0}\ket{l}\braket{l|\mathrm{i}[\rho,B]|k}\bra{k}$ (156)
$|\displaystyle=\mathrm{i}[\rho,B]-\sum_{k,l;p_{l}=p_{k}=0}\ket{l}\braket{l|\mathrm{i}[\rho,B]|k}\bra{k}$ (157)
@@
[…]
@@ B.2 Proof of the equivalence between two alternative definitions of asymptotic convertibility
[…]
@@ B.3 Proof of Lemma 4
Let $\sigma$ be an arbitrary state and $\phi$ be a pure state. Suppose that their infidelity by $\delta\coloneqq 1-\braket{\phi|\sigma|\phi}$ satisfies $\delta<1/2$. For the eigenvector $\Phi$ with the largest eigenvalue of $\sigma$, it holds
$|\braket{\phi|\Phi}|^{2}\geq 1-2\delta.$ (162)
@@
$\braket{\mathrm{i}[\sigma,O],\mathrm{i}[\sigma,O]}_{f,\sigma}\geq\frac{1}{f\left(\frac{\delta}{1-\delta}\right)}(1-2\delta)^{2}\Braket{\Phi|O(I-\Phi)O^{\dagger}|\Phi}.$ (163)
@@
[…]
$1-\delta=p|\braket{\phi|\Phi}|^{2}+\sum_{k=1}^{d-1}\lambda_{k}|\braket{\phi|k}|^{2}\leq p\sum_{k=0}^{d-1}|\braket{\phi|k}|^{2}=p.$ (165)
@@
$1-\delta$ $=\braket{\phi|\sigma|\phi}$ (166)
$=p|\braket{\phi|\Phi}|^{2}+\sum_{k=1}^{d-1}\lambda_{k}|\braket{\phi|k}|^{2}$ (167)
$\leq p|\braket{\phi|\Phi}|^{2}+(1-p)\sum_{k=1}^{d-1}|\braket{\phi|k}|^{2}$ (168)
$\leq p|\braket{\phi|\Phi}|^{2}+\delta(1-|\braket{\phi|\Phi}|^{2}),$ (169)
@@
$|\braket{\phi|\Phi}|^{2}\geq\frac{1-2\delta}{p-\delta}\geq 1-2\delta,$ (170)
@@
$\braket{\mathrm{i}[\sigma,O],\mathrm{i}[\sigma,O]}_{f,\sigma}$ $=\sum_{k,l;\lambda_{l}f(\lambda_{k}/\lambda_{l})>0}\frac{1}{\lambda_{l}f(\lambda_{k}/\lambda_{l})}\braket{k|[\sigma,O^{\dagger}]|l}\braket{l|[O,\sigma]|k}$ (171)
$=\sum_{k,l;\lambda_{l}f(\lambda_{k}/\lambda_{l})>0}\frac{(\lambda_{l}-\lambda_{k})^{2}}{\lambda_{l}f(\lambda_{k}/\lambda_{l})}|\braket{l|O|k}|^{2}+\sum_{k,l;\lambda_{l}f(\lambda_{k}/\lambda_{l})>0,l\neq 0}\frac{(\lambda_{l}-\lambda_{k})^{2}}{\lambda_{l}f(\lambda_{k}/\lambda_{l})}|\braket{l|O|k}|^{2}$ (172)
$\geq\sum_{k;pf(\lambda_{k}/p)>0}\frac{(p-\lambda_{k})^{2}}{pf(\lambda_{k}/p)}|\braket{\Phi|O|k}|^{2}$ (173)
$=\sum_{k=1}^{d-1}\frac{(p-\lambda_{k})^{2}}{pf(\lambda_{k}/p)}|\braket{\Phi|O|k}|^{2},$ (174)
@@
$\braket{\mathrm{i}[\sigma,O],\mathrm{i}[\sigma,O]}_{f,\sigma}$ $\geq\sum_{k=1}^{d-1}\frac{(p-\lambda_{k})^{2}}{pf(\lambda_{k}/p)}|\braket{\Phi|O|k}|^{2}$ (175)
$\geq\frac{1}{f\left(\frac{\delta}{1-\delta}\right)}(1-2\delta)^{2}\sum_{k=1}^{d-1}|\braket{\Phi|O|k}|^{2}$ (176)
$=\frac{1}{f\left(\frac{\delta}{1-\delta}\right)}(1-2\delta)^{2}\braket{\Phi|O(I-\Phi)O^{\dagger}|\Phi}.$ (177)
@@
$\Braket{\Phi_{N}|O_{N}(I-\Phi_{N})O_{N}^{\dagger}|\Phi_{N}}$ $=V(\Phi_{N},A_{N})+V(\Phi_{N},B_{N})-\mathrm{i}\braket{\Phi_{N}|[A_{N},B_{N}]|\Phi_{N}},$ (183)
where $V$ denotes the variance $V(\Phi_{N},A_{N})\coloneqq\braket{\Phi_{N}|A_{N}^{2}|\Phi_{N}}-\braket{\Phi_{N}|A_{N}|\Phi_{N}}^{2}$. We derive a lower bound for each contribution in Eq. (183) under the assumption that $\Phi_{N}$ is sufficiently close to an i.i.d. copies of a pure state $\phi^{\otimes N}$.
@@
$-\mathrm{i}\braket{\Phi_{N}|[A_{N},B_{N}]|\Phi_{N}}$ $\geq-\mathrm{i}\braket{\phi^{\otimes N}|[A_{N},B_{N}]|\phi^{\otimes N}}-N\|\mathrm{i}[A,B]\|_{\infty}\|\Phi_{N}-\phi^{\otimes N}\|_{1}.$ (190)
@@
[…]
@@
$\Delta\mu_{N}$ $\coloneqq\braket{\phi^{\otimes N}|A_{N}|\phi^{\otimes N}}-\braket{\Phi_{N}|A_{N}|\Phi_{N}},$ (192)
$Q_{N}$ $\coloneqq(A_{N}-\braket{\Phi_{N}|A_{N}|\Phi_{N}}I)^{2}=(A_{N}-\braket{\phi^{\otimes N}|A_{N}|\phi^{\otimes N}}I+\Delta\mu_{N}I)^{2}.$ (193)
@@
$\braket{\phi^{\otimes N}|(A_{N}-\braket{\Phi_{N}}{A_{N}|\Phi_{N}}I)^{2}|\phi^{\otimes N}|=}\braket{\phi^{\otimes N}|(A_{N}-\braket{\phi^{\otimes N}}{A_{N}|\phi^{\otimes N}}I)^{2}|\phi^{\otimes N}|+}|\Delta\mu_{N}|^{2},$ (194)
@@
Here, we derive an upper bound of $\mathrm{Tr}\left(Q_{N}(\phi^{\otimes N}-\Phi_{N})\right)$. Let us denote the eigenvalue decomposition of $A$ by $|A=\sum_{j=1}^{d}a_{j}\ket{j}\bra{j}$. Its i.i.d. extension is written as
[…]
@@
where $\mu\coloneqq\braket{\phi|A|\phi}$, $\sigma^{2}\coloneqq\braket{\phi|A^{2}|\phi}-\braket{\phi|A|\phi}^{2}$, and $x>0$ is a positive parameter that will be fixed later.
@@
$Q_{N}^{\mathrm{(core)}}(x)$ $|,\displaystyle\coloneqq\sum_{\bm{j}\in\mathcal{A}^{(\mathrm{core})}(x)}(s(\bm{j})-N\mu+\Delta\mu_{N})^{2}\ket{\bm{j}}\bra{\bm{j}}$, (199)
$Q_{N}^{\mathrm{(tail)}}(x)$ $|.\displaystyle\coloneqq\sum_{\bm{j}\in\mathcal{A}^{(\mathrm{tail})}(x)}(s(\bm{j})-N\mu+\Delta\mu_{N})^{2}\ket{\bm{j}}\bra{\bm{j}}$. (200)
@@
Let $p$ denote a probability distribution defined by $p(j)\coloneqq|\braket{j|\phi}|^{2}$. We denote its i.i.d. extension by
@@
For sufficiently small $\epsilon>0$ satisfying $1-2\epsilon-g_{0}(\epsilon)>0$, we have
@@
holds for all sufficiently large $N$, as long as $\epsilon>0$ is sufficiently small so that $1-2\epsilon-g_{0}(\epsilon)>0$. ∎
From Lemmas 25 and 26, we obtain the following corollary:
@@
[…]
$\Braket{\Phi_{N}|O_{N}(I-\Phi_{N})O_{N}^{\dagger}|\Phi_{N}}\geq\braket{\phi^{\otimes N}|O_{N}(I-\phi^{\otimes N})O_{N}^{\dagger}|\phi^{\otimes N}}-Nh_{2}(\epsilon)$ (229)
@@
$\Braket{\Phi_{N}|O_{N}(I-\Phi_{N})O_{N}^{\dagger}|\Phi_{N}}\geq\braket{\phi^{\otimes N}|O_{N}(I-\phi^{\otimes N})O_{N}^{\dagger}|\phi^{\otimes N}}-Nh_{2}(\epsilon)$ (231)
@@
$\braket{\mathrm{i}[\sigma_{N},O],\mathrm{i}[\sigma_{N},O]}_{f,\sigma_{N}}\geq\frac{1}{f\left(\frac{\delta_{N}}{1-\delta_{N}}\right)}(1-2\delta_{N})^{2}\Braket{\Phi_{N}|O_{N}(I-\Phi)O_{N}^{\dagger}|\Phi_{N}},$ (234)
where $\delta_{N}\coloneqq 1-\braket{\phi^{\otimes N}|\sigma_{N}|\phi^{\otimes N}}$ denotes the infidelity. For a sufficiently small $\epsilon>0$, from Lemma 27, we further find
$\braket{\mathrm{i}[\sigma_{N},O],\mathrm{i}[\sigma_{N},O]}_{f,\sigma_{N}}\geq\frac{1}{f\left(\frac{\delta_{N}}{1-\delta_{N}}\right)}(1-2\delta_{N})^{2}\left(N\braket{\phi|O(I-\phi)O^{\dagger}|\phi}-Nh_{2}(\epsilon)\right)$ (235)
@@
$f(0)\braket{\mathrm{i}[\sigma_{N},O],\mathrm{i}[\sigma_{N},O]}_{f,\sigma_{N}}\geq\frac{f(0)}{f\left(\frac{\delta_{N}}{1-\delta_{N}}\right)}(1-2\delta_{N})^{2}N\braket{\phi|O(I-\phi)O^{\dagger}|\phi}-Nh(\epsilon),$ (236)
@@
$f(0)\braket{\mathrm{i}[\sigma_{N},O],\mathrm{i}[\sigma_{N},O]}_{f,\sigma_{N}}\geq N\braket{\phi|O(I-\phi)O^{\dagger}|\phi}-Nh(\epsilon)+N\braket{\phi|O(I-\phi)O^{\dagger}|\phi}\left(\frac{f(0)}{f\left(\frac{\delta_{N}}{1-\delta_{N}}\right)}(1-2\delta_{N})^{2}-1\right).$ (238)